OpenStudy (anonymous):

Find the sum Y/y^2-2y+1 + 6/y^2+6y-7

OpenStudy (anonymous):

Do you want me to explain it to you after I solve it? What part of the question do you need help on?

OpenStudy (anonymous):

\[\frac{ y }{ y^2-2y+1 }+\frac{ 6 }{ y^2+6y-7 }\]Is this the question?

OpenStudy (anonymous):

Yes that's the question and please explain too

OpenStudy (anonymous):

Alright. Do you know how to factor?

OpenStudy (anonymous):

Yes

OpenStudy (rhr12):

take a look there

OpenStudy (anonymous):

Okay here are the steps then:\[\frac{ y }{ (y-1)^2 }+\frac{ 6 }{ (y-1)(y+7) }\]

OpenStudy (anonymous):

Do you know how I got to that?

OpenStudy (anonymous):

Are you still there?

OpenStudy (zale101):

Every time i deal with quotients adding together (even for fraction) i always abide by this rule: \(\LARGE \frac{A}{B}+\frac{C}{D}\) \(\Large\frac{A*D+C*B}{B*D}\)

OpenStudy (anonymous):

Next step:\[(y-1)(\frac{ y }{ (y-1)^2 }+\frac{ 6 }{ (y-1)(y+7) })\]

OpenStudy (anonymous):

\[\frac{ y }{ (y-1) }+\frac{ 6 }{ (y+7) }\]

OpenStudy (anonymous):

That would be the sum. You can still go further if you are trying to solve for y.